OpenStudy (anonymous):

Sofia joined a golf club two years ago. She pays an annual membership fee of $1,045 and a greens fee of $35 each time she plays a game of golf. The function below can be used to calculate the total yearly golfing fee, f (g), in dollars. f (g) = 1045 + 35g, where g represents the number of times she played golf during the year. Last year she paid $1,710 as a total golfing fee. For how many games did she pay a greens fee?

hero (hero):

f (g) = 1045 + 35g 1710 = 1045 + 35g 1710 - 1045 = 35g 665 = 35g 665/35 = g 19 = g

OpenStudy (anonymous):

Oh Okay now I see Thanks so much for your help hero :)

hero (hero):

? Mancake? What did you try?